OpenStudy (anonymous):

Write the equation of a line in point slope form that has a slope of -3/4 and passes through the point (4, 5).

OpenStudy (anonymous):

ok, so first you put it into slope intercept form, which is y=mx+b where you replace m with (-3,4), to get y=(-3/4)x, now for it to pass through the coordinates just add the x coordinate to the end twice, y=(-3/4)x+4+4, which simplified comes to what your looking for.

OpenStudy (anonymous):

so y=-3/4+4 is the answer

OpenStudy (anonymous):

y=(-3/4)x+4+4 .....................^ just add those and thats it

OpenStudy (anonymous):

or to put it another way y=-3/4x+(4+4) just do the parentheses and thats your answer

OpenStudy (anonymous):

(in slope intercept form)

OpenStudy (anonymous):

so y=-3/4+8

OpenStudy (anonymous):

y=-3/4x+8

OpenStudy (anonymous):

yes thats it

OpenStudy (anonymous):

thanks